OpenStudy (anonymous):

Liam is buying books for his summer reading collection. He has narrowed his choices down to 7 books and is planning to purchase 3. How many combinations exist?

OpenStudy (anonymous):

guess that would be \[\dbinom{7}{3}\]

OpenStudy (anonymous):

\[\frac{7\times 6\times 5}{3\times 2}=7\times 5=35\]